# A Survey on Bundle Recommendation: Methods, Applications, and Challenges

### v1

@@ Abstract.
copyright: nonecopyright: acmlicensedjournalyear: $0$doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 0journalnumber: 0article: 000publicationmonth: 0
@@ 1. Introduction
Recommendation systems have become essential tools for alleviating information overload. Over the years, item recommendation systems have been widely applied to recommend item to user or users across various domains, achieving significant success (Gomez-Uribe and Hunt, 2016; Das et al., 2007; Briand et al., 2021; Cheuque et al., 2019; Gao et al., 2023a). For example, in news recommendation, users’ reading history and interests are analyzed to recommend relevant articles, thereby increasing user engagement and satisfaction (Das et al., 2007; Liu et al., 2010; Phelan et al., 2009; Okura et al., 2017). In e-commerce, recommendation systems have become a cornerstone for boosting sales conversion rates by recommending products related to users’ browsing or purchase history, as evidenced by the sophisticated algorithms used by platforms like Amazon (McAuley et al., 2015a). Similarly, video streaming platforms like YouTube employ deep learning techniques to ensure users can discover videos or series they are interested in, significantly increasing watch time and user retention (Covington et al., 2016). Netflix’s recommendation system stands out as a prime example of leveraging complex algorithms and data analysis to provide highly personalized viewing recommendations, which has significantly enhanced user experience and subscription retention (Gomez-Uribe and Hunt, 2016). Music streaming services like Spotify use collaborative filtering (Schafer et al., 2007; Sarwar et al., 2001; Breese et al., 1998) and content-based methods (Lops et al., 2011; Pazzani and Billsus, 2007; Lu et al., 2015) to recommend individual tracks tailored to users’ preferences, improving user satisfaction and engagement (Koren et al., 2009; van den Oord et al., 2013). Additionally, social media platforms such as Facebook and Twitter utilize recommendation systems to deliver personalized content to users, thereby boosting user interaction and time spent on the platform (Guy et al., 2010; Kazai et al., 2016). Even in the travel and hospitality industry, recommendation systems help users find ideal destinations, accommodations, and activities based on their past behaviors and preferences (Linden et al., 2003).
@@
For illustration, several example bundles in the domain of Product, Clothing, Food, Entertainment, Health etc. are depicted in Figure 1. In e-commerce, Bundle RS assists retailers in offering product bundles tailored to specific customer needs and preferences, thereby enhancing the shopping experience and boosting sales (Liu et al., 2017, 2014; Wan et al., 2018; Zhu et al., 2014). For instance, when an electronics enthusiast is looking to buy an iPhone, Bundle RS might recommend complementary items such as AirPods, Apple Watch, and 3-in-1 charger. The electronic product bundles are often sold at a discount, further encouraging the purchase. Similarly, in fashion outfit recommendation (Liu et al., 2012; Han et al., 2017; Hu et al., 2015; McAuley et al., 2015b; Chen et al., 2019a; Li et al., 2020b; Lu et al., 2019), the fourth example of Figure 1 shows an outfit, including blouse, skirt, heels and a handbag. Suppose that a customer is shopping online with an intent, e.g., purchasing fashion clothing for a party, and finds this well-matched outfit appealing. In this case, the customer may purchase this outfit.
[…]
Different application domains of bundles
@@
In the entertainment industry, streaming services like Netflix, Spotify, and NetEase use Bundle RS to recommend sets of movies, TV shows, or songs, providing users with a more comprehensive and enjoyable viewing or listening experience (Gomez-Uribe and Hunt, 2016; Cao et al., 2017; Chang et al., 2020; He et al., 2020b; Yang et al., 2019). Similarly, platforms like Youshu and Goodreads provide lists of books created by readers, leveraging Bundle RS to offer curated book recommendations (Chen et al., 2019b; He et al., 2019b; Liu et al., 2014). These bundles enhance the reading experience by recommending books that complement each other, whether based on themes, genres, or reader preferences. Meal recommendation (Agapito et al., 2017; Elsweiler and Harvey, 2015; Li et al., 2022a, 2023b, 2023a) is another application of bundle recommendation in food scenarios. For instance, a user looking for a dinner idea might receive a recommendation for a three-course meal featuring a broccoli salad, pasta, and an ice cream cake for dessert. This holistic approach to meal planning simplifies decision-making and can inspire users to try new recipes and ingredients. In the realm of travel package recommendation (Herzog et al., 2019; Lim et al., 2015, 2018), suppose that a family is planning a vacation to an unfamiliar city, Bundle RS will recommend a travel package that includes an ordered set of Point of Interest (POIs) that align with their interest preferences and trip constraints. This makes trip planning more convenient and ensures a memorable and well-rounded travel experience. Since bundles are ubiquitous across various application domains, Bundle RS is widely used in e-commerce, entertainment, and travel industries and so on.
@@
Bundle RS has become crucial in enhancing user experiences and boosting business success. In a world saturated with choices, users often feel overwhelmed by the need to make decisions. By offering personalized bundles of items, such as book lists, song lists, meals, or travel package, Bundle RS simplifies decision-making and provides cohesive, enjoyable experiences for users. This task not only increases user satisfaction and loyalty but also drives higher sales, as seen in e-commerce where tailored product bundles are used to meet specific user needs and preferences.
@@
With the growing demand and development of bundle recommendation across various domains, numerous studies have been conducted to improve its performance. Initially, Zhu et al. (Zhu et al., 2014) proposed the bundle recommendation problem and tried to resolve it by minimizing the cost or maximize the revenue of a bundle (Zhu et al., 2014; Beladev et al., 2016). Later, association rule mining was utilized in (Guo-rong and Xi-zheng, 2006; Fang et al., 2018) for bundle generation and recommendation. When a certain number of user-bundle interactions were available, an intuitive solution was to treat the bundle as a single 'item' and apply traditional collaborative filtering methods. Prior work (Rendle et al., 2010) just ignored the affiliated items of the bundle and just used an id to represent a bundle. Recognizing the importance of affiliated items, some works (Cao et al., 2017; Chen et al., 2019b) tried to capture additional user-item interaction and bundle-item affiliation relations. Over time, the bundle view and item view were gradually introduced in bundle recommendation research (Chang et al., 2020; Ma et al., 2022) etc. This shift led to an increasing amount of research focused on these two views within bundle recommendation (Ke et al., 2023b, a; Ren et al., 2023) etc. Considering the increasing attention, emerging challenges, and urgent need for innovative breakthroughs on bundle recommendation, we realize it is the right time to present a survey of this area, providing a systematic review of the methods, applications, and challenges in Bundle RS.
@@
Given the significance and popularity of recommendation research, there are more and more surveys reviewing RS from different perspectives. These include explainable recommendation (Zhang and Chen, 2020), knowledge-based recommendation (Tarus et al., 2018), recommendation methods based on deep learning and reinforcement learning (Chen et al., 2023b), accuracy-oriented recommendation modeling (Wu et al., 2023b), bias and debias in RS (Chen et al., 2023a), fairness in RS (Deldjoo et al., 2024; Wang et al., 2023b), and so on. A survey on bundle recommendation by (Li et al., 2020a) published in 2020 primarily reviews the main model methods of bundle RS, including integer programming, association analysis, and both traditional and deep learning-based recommendation techniques. These methods have provided effective solutions in specific application scenarios. Along with the development of new techniques such as graph learning (Chang et al., 2020; Li et al., 2023b; Wei et al., 2023), contrastive learning (Zhao et al., 2022; Ma et al., 2024a, 2022), and knowledge distillation (Ren et al., 2023), the research on bundle recommendation has made further progress in the past few years. Therefore, different from existing surveys, our survey offers a more comprehensive and up-to-date review on bundle recommendation, covering the datasets across various bundle recommendation scenarios, both discriminative and generative bundle recommendation tasks, representation learning approaches and strategies, as well as interaction prediction and bundle generation methods. This survey benefits for researchers and practitioners who want to keep up with the state-of-the-art research in the field of bundle recommendation. By doing so, we deliver the following contributions in this survey:
@@
[…]
The statistics and distribution of papers related to bundle recommendation
@@
This survey is structured into seven sections and is organized as follows: Section 2 introduces the definitions and taxonomy of bundle recommendation. Section 3 discusses the discriminative bundle recommendation task, reviews the representation learning techniques and strategies from bundle and item levels, and summarizes several approaches of interaction modeling. Section 4 explores the generative bundle recommendation task, and reviews representation learning from item level and bundle generation methods. Section 5 summarizes the resources of bundle recommendation and conducts reproducibility experiments. Section 6 discusses the challenges and trends. And Section 7 concludes the survey.
@@ 2. Definitions and Taxonomy
In this section, we introduce the lifecycle of bundle recommendation, providing an overview of its key stages. Based on it, we define the necessary notations, and formulate the problem and task of bundle recommendation. Finally, we categorize bundle recommendation into two distinct types based on bundle composition strategy.
@@ 2.2. Problem Formulation
Let’s use $\mathcal{U}=\{u_{1},u_{2},...,u_{M}\}$, =$\mathcal{B}=\{b_{1},b_{2},...,b_{N}\}$, and =$\mathcal{I}=\{i_{1},i_{2},...,i_{O}\}$ to denote the sets of users, bundles and items. $M$, $N$ and $O$ denote the number of users, bundles and items, respectively. Each bundle $b=\{i_{1},i_{2},...,i_{K}\}$, $b\in\mathcal{B}$, $i\in\mathcal{I}$. $K$ is the size of bundle $b$.
@@ 2.3. Taxonomy
[…]
Taxonomy of bundle recommendation
Figure 4. Taxonomy of Bundle Recommendation
Figure 4 illustrates the taxonomy of bundle recommendation, discriminative and generative, respectively. It outlines the respective tasks for different categories and proposes the corresponding solutions. Datasets and evaluation metrics are the application resources in bundle recommendation. The details are presented in the following sections.
@@ 3.2. Representation Learning
This part is aim to learn high-quality representations of users, items, and bundles from bundle and item levels based on different learning strategies. As illustrated in Table 2, we categorize different models according to representation learning approaches and strategies. The representation approach includes factorization models (Cao et al., 2017; Liu et al., 2014; Chen et al., 2019b; Pathak et al., 2017), attention-based models (He et al., 2019b; Li et al., 2021c; Yang et al., 2020; Ren et al., 2023; Vijaikumar et al., 2021), and GNNs-based models (Chang et al., 2020; Deng et al., 2020; Ma et al., 2022; Li et al., 2023b; Ma et al., 2024a). We will introduce the different approaches in Section 3.2.1 specifically. As for learning strategy, we divide them into the following three types, namely unified representation learning, separate representation learning, and cooperative representation learning. This part will be discussed in Section 3.2.2.
@@
BundleNet(Deng et al., 2020), PGAT(Lu et al., 2022), CLBR(Zhu et al., 2023c)
HED(Liu et al., 2023a), UHBR(Yu et al., 2022), HGLR(Yang et al., 2023)
HFGN(Li et al., 2020b), AGF(Li et al., 2023c)
@@
DT-CDBR(Zhang et al., 2021b), CMRec(Li et al., 2023b), TGIF(Kim et al., 2024)
@@
Cooperative - DAM(Chen et al., 2019b), BundleBPR(Pathak et al., 2017) GANCF(Yang et al., 2020), DAM(Chen et al., 2019b) IMBR(Sun et al., 2022a)
@@
CL2BRec(Wang et al., 2023c), CoHEAT(Jeon et al., 2024), EBRec(Du et al., 2023)
@@ 3.2.1. Backbone Models of Representation Learning
Previous research (Rendle et al., 2010; Min et al., 2020) did not consider bundles as collections of items; instead, it treated bundles as atomic units, similar to individual items. Then MF-based collaborative filtering is employed to learn bundle-level user preferences from user-bundle interaction matrix. Later, the distinction between bundles and items has received increasing attention. With the development of pairwise ranking approach, some works have attempted to rank the bundle by optimizing a pairwise ranking loss. LIRE (Liu et al., 2014), EFM (Cao et al., 2017) and BundleBPR (Pathak et al., 2017) simultaneously utilize the users’ interactions with both items and bundles under the BPR framework (Rendle et al., 2009). The bundle vectors can be aggregated from the latent vectors of the items they contain. By optimizing these different latent factor vectors using the Bayesian Personalized Ranking (BPR) loss, factorization models can learn comprehensive and effective representations of both users and bundles. Furthermore, BundleBPR model (Pathak et al., 2017) leverages the parameters learned through the item BPR model to estimate the preference of a user toward a bundle. Recognizing the rich signal about item co-occurrence in bundles, ProductRec (Yu et al., 2017) explores the latent bundling relationships between items based on FPMC. EFM (Cao et al., 2017) believes that discovering the relationship between bundles and their contained items can mutually reinforce the performance of user-item and user-bundle recommendation. Out of this thought, they devise embedding factorization models, which jointly factorize user-item interaction matrix, user-bundle interaction matrix, and item-item-bundle co-occurrence matrix within a unified framework. Combining neural network techniques, DAM (Chen et al., 2019b) proposes a factorized attention network, which factorizes the attention weight matrix with a low-rank model followed by a softmax operation, differing from existing designs of neural attention network. In particular, considering the heterogeneous features of items in a bundle, FPITF (Hu et al., 2015) proposes a functional tensor factorization model for fashion recommendation, which decomposes the high order interactions between users and the fashion items into a set of pairwise interactions in some latent space.
@@
However, the more items two bundles have in common, the more similar they are, since a bundle is composed of multiple items. Therefore, it is inappropriate to treat bundles as separate columns in the user-bundle matrix to run CF models like matrix factorization. To address the compositional similarity between bundles, it is essential to aggregate item embeddings for bundle representation. Since items within a bundle have varying degrees of importance, the attention mechanism is employed to dynamically weight the importance of different items during aggregation (Chen et al., 2019b; He et al., 2019b; Li et al., 2021c; Yang et al., 2023). For each item $i$ in the bundle $b$, an attention score $e_{bi}$ is calculated. The attention score is computed using a combination of item and user embeddings:
@@
(7) $\mathbf{b}_{b}=\sum_{i\in b}\alpha_{bi}\mathbf{v}_{i}.$
One representative work is DAM (Chen et al., 2019b) which designs a factorized attention network to aggregate the item embeddings in a bundle to obtain the bundle’s representation. The design of the weighting strategy can be seen as factorizing the attention weight matrix with a low-rank model. Besides the intra-connections among the items within the bundle, there are also inter-connections among bundles. IPRec (Li et al., 2021c) considers the inter-connection relationships for bundle modeling, and proposes a intra- and inter-bundle attention network for bundle recommendation. Similarly, AttList (He et al., 2019b) models latent representations of bundles and users via item-level and bundle-level attention networks. On the other hand, the items within a bundle are usually gathered according to a specific theme, which may contain some latent attributes. GANCF (Yang et al., 2020) proposes item-level and attribute-level attention networks to reinforce the representations of bundles. HGLR (Yang et al., 2023) introduces an attribute-based heterogeneous information network for representation learning and utilizes an attention network to fuse the embeddings learned from bundle and item views.
@@
Based on users’ interactions with items and bundles, as well as affiliations between bundles and items, it is then necessary to connect user nodes with item nodes and bundle nodes. Once the graph has been constructed, the initial the embedding vector for each node (users, items, and bundles) should be randomly initialized or derived from pre-trained models. Subsequently, the message passing mechanism, a core operation in GNNs, should be implemented. This involves performing message passing between neighboring nodes to aggregate information from their local neighborhoods. After that, the representation of each node is updated based on the aggregated information. This update is typically performed using a neural network layer, such as a graph convolutional layer, which integrates information from neighboring nodes into the node’s representation. The update rule for a node $i$ at layer $k+1$ can be expressed as:

### v2

@@ Abstract.
journal: JACMccs: Information systems Recommender systems
@@ 1. Introduction
Recommendation systems (RS) have become essential tools for alleviating information overload. Over the years, item recommendation systems have been widely applied to recommend item to user or users across various domains, achieving significant success (Gomez-Uribe and Hunt, 2016; Das et al., 2007; Briand et al., 2021; Cheuque et al., 2019; Gao et al., 2023a). For example, in news recommendation, users’ reading history and interests are analyzed to recommend relevant articles, thereby increasing user engagement and satisfaction (Das et al., 2007; Liu et al., 2010; Phelan et al., 2009; Okura et al., 2017). In e-commerce, recommendation systems have become a cornerstone for boosting sales conversion rates by recommending products related to users’ browsing or purchase history, as evidenced by the sophisticated algorithms used by platforms like Amazon (McAuley et al., 2015a). Similarly, video streaming platforms like YouTube employ deep learning techniques to ensure users can discover videos or series they are interested in, significantly increasing watch time and user retention (Covington et al., 2016). Netflix’s recommendation system stands out as a prime example of leveraging complex algorithms and data analysis to provide highly personalized viewing recommendations, which has significantly enhanced user experience and subscription retention (Gomez-Uribe and Hunt, 2016). Music streaming services like Spotify use collaborative filtering (Schafer et al., 2007; Sarwar et al., 2001; Breese et al., 1998) and content-based methods (Lops et al., 2011; Pazzani and Billsus, 2007; Lu et al., 2015) to recommend individual tracks tailored to users’ preferences, improving user satisfaction and engagement (Koren et al., 2009; van den Oord et al., 2013). Additionally, social media platforms such as Facebook and Twitter utilize recommendation systems to deliver personalized content to users, thereby boosting user interaction and time spent on the platform (Guy et al., 2010; Kazai et al., 2016). Even in the travel and hospitality industry, recommendation systems help users find ideal destinations, accommodations, and activities based on their past behaviors and preferences (Linden et al., 2003).
@@
For illustration, several example bundles in the domain of Product, Clothing, Food, Entertainment, Health etc. are depicted in Figure 1. In e-commerce, Bundle RS assists retailers in offering product bundles tailored to specific customer needs and preferences, thereby enhancing the shopping experience and boosting sales (Liu et al., 2017, 2014; Wan et al., 2018; Zhu et al., 2014). For instance, when an electronics enthusiast is looking to buy an iPhone, Bundle RS might recommend complementary items such as AirPods, Apple Watch, and 3-in-1 charger. The electronic product bundles are often sold at a discount, further encouraging the purchase. Similarly, in fashion outfit recommendation (Liu et al., 2012; Han et al., 2017; Hu et al., 2015; McAuley et al., 2015b; Chen et al., 2019b; Li et al., 2020b; Lu et al., 2019), the fourth example of Figure 1 shows an outfit, including blouse, skirt, heels and a handbag. Suppose that a customer is shopping online with an intent, e.g., purchasing fashion clothing for a party, and finds this well-matched outfit appealing. In this case, the customer may purchase this outfit.
@@
In the entertainment industry, streaming services like Netflix, Spotify, and NetEase use Bundle RS to recommend sets of movies, TV shows, or songs, providing users with a more comprehensive and enjoyable viewing or listening experience (Gomez-Uribe and Hunt, 2016; Cao et al., 2017; Chang et al., 2020; He et al., 2020b; Yang et al., 2019). Similarly, platforms like Youshu and Goodreads provide lists of books created by readers, leveraging Bundle RS to offer curated book recommendations (Chen et al., 2019a; He et al., 2019b; Liu et al., 2014). These bundles enhance the reading experience by recommending books that complement each other, whether based on themes, genres, or reader preferences. Meal recommendation (Agapito et al., 2017; Elsweiler and Harvey, 2015; Li et al., 2022b, 2023a, 2023b) is another application of bundle recommendation in food scenarios. For instance, a user looking for a dinner idea might receive a recommendation for a three-course meal featuring a broccoli salad, pasta, and an ice cream cake for dessert. This holistic approach to meal planning simplifies decision-making and can inspire users to try new recipes and ingredients. In travel package recommendation (Herzog et al., 2019; Lim et al., 2015, 2018), suppose a family is planning a vacation to an unfamiliar city; Bundle RS will recommend a travel package with an ordered set of Points of Interest (POIs) that align with their preferences and trip constraints. This makes trip planning more convenient and ensures a memorable, well-rounded experience. Since bundles are ubiquitous across application domains, Bundle RS is widely used in e-commerce, entertainment, and travel.
@@
Bundle RS has become crucial in enhancing user experiences and boosting business success. In a world saturated with choices, users often feel overwhelmed by decisions. By offering personalized bundles, such as book lists, song lists, meals, or travel package, Bundle RS simplifies decision-making and provides cohesive, enjoyable experiences for users. This task not only increases user satisfaction and loyalty but also drives higher sales, as seen in e-commerce where tailored product bundles are used to meet specific user needs and preferences.
@@
With the growing demand and development of bundle recommendation across various domains, numerous studies have been conducted to improve its performance. Initially, Zhu et al. (Zhu et al., 2014) proposed the bundle recommendation problem and tried to resolve it by minimizing the cost or maximize the revenue of a bundle (Zhu et al., 2014; Beladev et al., 2016). Later, association rule mining was utilized in (Guo-rong and Xi-zheng, 2006; Fang et al., 2018) for bundle generation and recommendation. When a certain number of user-bundle interactions were available, an intuitive solution was to treat the bundle as a single 'item' and apply traditional collaborative filtering methods. Prior work (Rendle et al., 2010) just ignored the affiliated items of the bundle and just used an id to represent a bundle. Recognizing the importance of affiliated items, some works (Cao et al., 2017; Chen et al., 2019a) tried to capture additional user-item interaction and bundle-item affiliation relations. Over time, the bundle view and item view were gradually introduced in bundle recommendation research (Chang et al., 2020; Ma et al., 2022) etc. This shift led to an increasing amount of research focused on these two views within bundle recommendation (Ke et al., 2023b, a; Ren et al., 2023) etc. Considering the increasing attention, emerging challenges, and urgent need for innovative breakthroughs on bundle recommendation, we realize it is the right time to present a survey of this area, providing a systematic review of the methods, applications, and challenges in Bundle RS.
@@
Given the significance and popularity of recommendation research, there are more and more surveys reviewing RS from different perspectives. These include explainable recommendation (Zhang and Chen, 2020), knowledge-based recommendation (Tarus et al., 2018), recommendation methods based on deep learning and reinforcement learning (Chen et al., 2023b), accuracy-oriented recommendation modeling (Wu et al., 2023a), bias and debias in RS (Chen et al., 2023a), fairness in RS (Deldjoo et al., 2024; Wang et al., 2023c), and so on. A survey on bundle recommendation by (Li et al., 2020a) published in 2020 primarily reviews the main model methods of bundle RS, including integer programming, association analysis, and both traditional and deep learning-based recommendation techniques. These methods have provided effective solutions in specific application scenarios. Along with the development of new techniques such as graph learning (Chang et al., 2020; Li et al., 2023a; Wei et al., 2023), contrastive learning (Zhao et al., 2022; Ma et al., 2024a, 2022), and knowledge distillation (Ren et al., 2023), the research on bundle recommendation has made further progress in the past few years. Therefore, different from existing surveys, our survey offers a more comprehensive and up-to-date review on bundle recommendation, covering the datasets across various bundle recommendation scenarios, both discriminative and generative bundle recommendation tasks, representation learning approaches and strategies, as well as interaction prediction and bundle generation methods. This survey benefits for researchers and practitioners who want to keep up with the state-of-the-art research in the field of bundle recommendation. By doing so, we deliver the following contributions in this survey:
@@
This survey is structured into seven sections and is organized as follows: Section 2 introduces the definitions and taxonomy of bundle recommendation. Section 3 discusses the discriminative bundle recommendation task, reviews the representation learning techniques and strategies from bundle and item levels, and summarizes several approaches of interaction modeling. Section 4 explores the generative bundle recommendation task, and reviews representation learning from item level and bundle generation methods. Section 5 summarizes the resources of bundle recommendation. Section 6 discusses the challenges and trends. And Section 7 concludes the survey.
@@ 2. Definitions and Taxonomy
In this section, we introduce the lifecycle of bundle recommendation, providing an overview of its key stages. Based on it, we define the necessary notations, formulate the problem and task of bundle recommendation, and summarize the abbreviations used in Table 2 for clarity. Finally, we categorize bundle recommendation into two distinct types based on bundle composition strategy.
Table 2. Full Names and Abbreviations
Full Name Abbreviation Full Name Abbreviation
Recommender System RS User-Bundle Interaction U-B
Bundle Recommendation System Bundle RS User-Item Interaction U-I
Discriminative Bundle Recommendation Discriminative BR Bundle-Item Interaction B-I
Generative Bundle Recommendation Generative BR Knowledge Distillation KD
Graph Neural Network GNN Point of Interest POI
Convolutional Neural Network CNN Latent Dirichlet Allocation LDA
Recurrent Neural Network RNN Term Transition $T$
Generative Adversarial Network GAN Determinantal Point Process DPP
Long Short-Term Memory LSTM Natural Language Processing NLP
Collaborative Filtering CF Large Language Model LLM
@@ 2.2. Problem Formulation
Let’s use $\mathcal{U}=\{u_{1},u_{2},...,u_{M}\}$, $\mathcal{B}=\{b_{1},b_{2},...,b_{N}\}$, and $\mathcal{I}=\{i_{1},i_{2},...,i_{O}\}$ to denote the sets of users, bundles and items. $M$, NN and $O$ denote the number of users, bundles and items, respectively. Each bundle $\{i_{1},i_{2},...,i_{K}\}, bb\in\mathcal{B}$, $i\in\mathcal{I}$. $K$ is the size of bundle $b$.
@@ 2.3. Taxonomy
Figure 4 illustrates the taxonomy of bundle recommendation, discriminative and generative, respectively. It outlines the respective tasks for different categories and proposes the corresponding solutions. Datasets and evaluation metrics are the application resources in bundle recommendation. The details are presented in the following sections. It is worth noting that this survey primarily focuses on general bundle recommendation tasks. Domain-specific scenarios (e.g., drug package recommendation, where requires detailed pharmacological knowledge) are only briefly mentioned but not discussed in depth, as they rely on strong domain expertise beyond the general scope of bundle recommendation systems.
@@ 3.2. Representation Learning
This part is aim to learn high-quality representations of users, items, and bundles from bundle and item levels based on different learning strategies. As illustrated in Table 3, we categorize different models according to representation learning approaches and strategies. The representation approach includes factorization models (Cao et al., 2017; Liu et al., 2014; Chen et al., 2019a; Pathak et al., 2017), attention-based models (He et al., 2019b; Li et al., 2021c; Yang et al., 2020; Ren et al., 2023; Vijaikumar et al., 2021), and GNNs-based models (Chang et al., 2020; Deng et al., 2020; Ma et al., 2022; Li et al., 2023a; Ma et al., 2024a). We will introduce the different approaches in Section 3.2.1 specifically. As for learning strategy, we divide them into the following three types, namely unified representation learning, separate representation learning, and cooperative representation learning. This part will be discussed in Section 3.2.2.
@@
BundleNet(Deng et al., 2020), PGAT(Lu et al., 2022), CLBR(Zhu et al., 2023b)
HED(Liu et al., 2024a), UHBR(Yu et al., 2022), HGLR(Yang et al., 2023)
HFGN(Li et al., 2020b), AGF(Li et al., 2023d)
@@
DT-CDBR(Zhang et al., 2021b), CMRec(Li et al., 2023a), TGIF(Kim et al., 2024)
@@
Cooperative - DAM(Chen et al., 2019a), BundleBPR(Pathak et al., 2017) GANCF(Yang et al., 2020), DAM(Chen et al., 2019a) IMBR(Sun et al., 2022a)
@@
CL2BRec(Wang et al., 2023a), CoHEAT(Jeon et al., 2024), EBRec(Du et al., 2023)
@@ 3.2.1. Backbone Models of Representation Learning
Previous research (Rendle et al., 2010; Min et al., 2020) did not consider bundles as collections of items; instead, it treated bundles as atomic units, similar to individual items. Then MF-based collaborative filtering is employed to learn bundle-level user preferences from user-bundle interaction matrix. Later, the distinction between bundles and items has received increasing attention. With the development of pairwise ranking approach, some works have attempted to rank the bundle by optimizing a pairwise ranking loss. LIRE (Liu et al., 2014), EFM (Cao et al., 2017) and BundleBPR (Pathak et al., 2017) simultaneously utilize the users’ interactions with both items and bundles under the BPR framework (Rendle et al., 2009). The bundle vectors can be aggregated from the latent vectors of the items they contain. By optimizing these different latent factor vectors using the Bayesian Personalized Ranking (BPR) loss, factorization models can learn comprehensive and effective representations of both users and bundles. Furthermore, BundleBPR model (Pathak et al., 2017) leverages the parameters learned through the item BPR model to estimate the preference of a user toward a bundle. Recognizing the rich signal about item co-occurrence in bundles, ProductRec (Yu et al., 2017) explores the latent bundling relationships between items based on FPMC. EFM (Cao et al., 2017) believes that discovering the relationship between bundles and their contained items can mutually reinforce the performance of user-item and user-bundle recommendation. Out of this thought, they devise embedding factorization models, which jointly factorize user-item interaction matrix, user-bundle interaction matrix, and item-item-bundle co-occurrence matrix within a unified framework. Combining neural network techniques, DAM (Chen et al., 2019a) proposes a factorized attention network, which factorizes the attention weight matrix with a low-rank model followed by a softmax operation, differing from existing designs of neural attention network. In particular, considering the heterogeneous features of items in a bundle, FPITF (Hu et al., 2015) proposes a functional tensor factorization model for fashion recommendation, which decomposes the high order interactions between users and the fashion items into a set of pairwise interactions in some latent space.
@@
However, the more items two bundles have in common, the more similar they are, since a bundle is composed of multiple items. Therefore, it is inappropriate to treat bundles as separate columns in the user-bundle matrix to run CF models like matrix factorization. To address the compositional similarity between bundles, it is essential to aggregate item embeddings for bundle representation. Since items within a bundle have varying degrees of importance, the attention mechanism is employed to dynamically weight the importance of different items during aggregation (Chen et al., 2019a; He et al., 2019b; Li et al., 2021c; Yang et al., 2023). For each item $i$ in the bundle $b$, an attention score $e_{bi}$ is calculated. The attention score is computed using a combination of item and user embeddings:
@@
(7) $\mathbf{b}_{b}=\textstyle\sum_{i\in b}\alpha_{bi}\mathbf{v}_{i}.$
One representative work is DAM (Chen et al., 2019a) which designs a factorized attention network to aggregate the item embeddings in a bundle to obtain the bundle’s representation. The design of the weighting strategy can be seen as factorizing the attention weight matrix with a low-rank model. Besides the intra-connections among the items within the bundle, there are also inter-connections among bundles. IPRec (Li et al., 2021c) considers the inter-connection relationships for bundle modeling, and proposes a intra- and inter-bundle attention network for bundle recommendation. Similarly, AttList (He et al., 2019b) models latent representations of bundles and users via item-level and bundle-level attention networks. On the other hand, the items within a bundle are usually gathered according to a specific theme, which may contain some latent attributes. GANCF (Yang et al., 2020) proposes item-level and attribute-level attention networks to reinforce the representations of bundles. HGLR (Yang et al., 2023) introduces an attribute-based heterogeneous information network for representation learning and utilizes an attention network to fuse the embeddings learned from bundle and item views.
@@
Bundle recommendation involves modeling interactions among users, items, and bundles. To capture these relations, a graph is constructed where nodes represent users, items, and bundles, and edges encode user–item interactions, user–bundle interactions, or bundle–item affiliations. Each node is initialized with an embedding vector, either randomly or from pre-trained models. A graph neural network (GNN) then updates node embeddings via message passing, aggregating information from neighbors. Formally, the update for node $i$ at layer $k+1$ is: